# Stock Price Prediction Using Triple Barrier Labeling and Raw OHLCV Data: Evidence from Korean Markets

### v1

@@ Abstract
This paper demonstrates that deep learning models trained on raw OHLCV (open-high-low-close-volume) data can achieve comparable performance to traditional machine learning models using technical indicators for stock price prediction in Korean markets. While previous studies have emphasized the importance of technical indicators and feature engineering, we show that a simple LSTM network trained on raw OHLCV data alone can match the performance of sophisticated ML models that incorporate technical indicators. Using a dataset of Korean stocks from 2006 to 2024, we optimize the triple barrier labeling parameters to achieve balanced label proportions with a 29-day window and 9% barriers. Our experiments reveal that LSTM networks achieve similar performance to traditional machine learning models like XGBoost, despite using only raw OHLCV data without any technical indicators. Furthermore, we identify that the optimal window size varies with model hidden size, with a configuration of window size 100 and hidden size 8 yielding the best performance. Additionally, our results confirm that using full OHLCV data provides better predictive accuracy compared to using only close price or close price with volume. These findings challenge conventional approaches to feature engineering in financial forecasting and suggest that simpler approaches focusing on raw data and appropriate model selection may be more effective than complex feature engineering strategies.
@@ III-A Data Collection and Preprocessing
    • Testing: September 29, 2022 – January 1, 2025
The resulting dataset contained a total of 8,566,617 instances for window size 100, with each instance having a shape of (5, 100), representing OHLCV sequences of length 100. Note that this dataset size is specific to a window size of 100, and may vary by approximately +/- 200 instances for other window settings.
@@ III-B Labeling with Triple Barrier Method
We employed the triple barrier labeling method to generate target labels for our classification task. This method defines labels based on the movement of a stock’s price in relation to three key barriers: a take-profit level, a stop-loss level, and a time limit. If the stock price reaches the take-profit or stop-loss level, it gets labeled accordingly. If neither barrier is hit within the specified time frame, the label reflects that no significant movement occurred. Key details are as follows:
@@ IV-C Feature Selection Analysis
  • Close price only (C) resulted in an F1 score of 0.2749
  • Adding volume (CV) increased the F1 score to 0.4251
@@
These results demonstrate that volume plays a critical role in prediction, while other OHLCV features provide incremental benefits when used together.
@@ IV-E Comparison with Previous Work
  • Raw OHLCV data can provide on par predictive power compared to engineered technical indicators when used with appropriately optimized deep learning models.
@@ V-A4 Feature Importance in OHLCV Data
Our feature selection experiments demonstrated that while close price alone is insufficient for accurate prediction (F1: 0.2749), adding volume substantially improves performance (F1: 0.4251), with the full OHLCV set providing further modest improvements (F1: 0.4312). This indicates that:
@@
    Volume as Key Indicator: Trading volume carries significant predictive information, possibly reflecting market sentiment and liquidity conditions not captured by price alone.
@@ V-B Conclusion
Finally, our results confirmed that using full OHLCV data provides better predictive accuracy compared to using only close price or close price with volume, with volume being particularly important for improving model performance.

### v2

@@ Abstract
This paper demonstrates that deep learning models trained on raw OHLCV (open-high-low-close-volume) data can achieve comparable performance to traditional machine learning (ML) models using technical indicators for stock price prediction in Korean markets. While previous studies have emphasized the importance of technical indicators and feature engineering, we show that a simple LSTM network trained on raw OHLCV data alone can match the performance of sophisticated ML models that incorporate technical indicators. Using a dataset of Korean stocks from 2006 to 2024, we optimize the triple barrier labeling parameters to achieve balanced label proportions with a 29-day window and 9% barriers. Our experiments reveal that LSTM networks achieve similar performance to traditional machine learning models like XGBoost, despite using only raw OHLCV data without any technical indicators. Furthermore, we identify that the optimal window size varies with model hidden size, with a configuration of window size 100 and hidden size 8 yielding the best performance. Additionally, our results confirm that using full OHLCV data provides better predictive accuracy compared to using only close price or close price with volume. These findings challenge conventional approaches to feature engineering in financial forecasting and suggest that simpler approaches focusing on raw data and appropriate model selection may be more effective than complex feature engineering strategies.
@@ III-A Data Collection and Preprocessing
    • Testing: September 29, 2022 – December 31, 2024
The resulting dataset contained a total of 8,566,617 instances for a window size of 100, with each instance having a shape of (5, 100), representing OHLCV sequences of length 100. For other window sizes, the instance count varies slightly (e.g., 8,566,522–8,566,817) due to differences in the number of valid rolling windows.
@@ III-B Labeling with Triple Barrier Method
We employed the triple barrier labeling method to generate target labels for our classification task. This method defines labels based on the movement of a stock’s price in relation to three key barriers: a take-profit level, a stop-loss level, and a time limit. If the stock price reaches the take-profit or stop-loss level, it gets labeled accordingly. If neither barrier is reached within the specified time frame, the label reflects that no significant movement occurred. Key details are as follows:
@@ IV-C Feature Selection Analysis
  • Close price only (C) resulted in an F1 score of 0.4170
  • Adding volume (CV) increased the F1 score to 0.4297
@@
The findings indicate that adding volume data improves the F1 score by 1.27%, while the remaining OHLCV features (open, high, low) together contribute only minimal additional performance gains.
@@ IV-E Comparison with Previous Work
  • Raw OHLCV data can provide comparable predictive power compared to engineered technical indicators when used with appropriately optimized deep learning models.
@@ V-A4 Feature Importance in OHLCV Data
Our feature selection experiments demonstrated that while close price alone provides substantial predictive power (F1: 0.4170), adding volume further improves performance (F1: 0.4297), with the full OHLCV set providing additional modest improvements (F1: 0.4312). This indicates that:
@@
    Volume as Additional Indicator: Trading volume provides supplementary predictive information, possibly reflecting market sentiment and liquidity conditions not captured by price alone.
@@ V-B Conclusion
Finally, our results confirmed that using full OHLCV data provides better predictive accuracy compared to using only close price or close price with volume, though each additional feature offers incremental improvements in model performance.